Find an equation in slope-intercept form for the line passing through each pair of points: (-4, 4), (-5, -3)
Answer by Alan3354(69443)

Find an equation in slope-intercept form for the line passing through each pair of points: (-4,4), (-5,-3)
Don't put a space after the comma.
-----------------
Do it like this one:
=============================
I need to find the equation of the line that passes through each pai of points and write the answer in standard form
(-2,-4), (1,-1)
--------------
Find the slope, m
m = diffy/diffx = (-4 + 1)/(-2 -1) = -3/-3 = 1
-----
Use y = mx + b and either point to find b
-1 = 1*1 + b
b = -2
-----
y = x - 2
x - y = 2
==================
Using determinants:
|+x +y 1|
|-2 -4 1| = 0
|+1 -1 1|
-------
x*(-4+1) - y*(-2-1) + 1*(2+4) = 0
-3x + 3y = -6
x - y = 2